# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Acknowledgements.
For modes with large-enough $p_{z}$, such that $p^{*}=4|p_{z}|\Omega/{U^{\prime}}>k_{f}$, and for ${U^{\prime}}/\Omega\lesssim 1$, forcing occurs in Sector H and waves there generate a decaying forward energy flux, $\Pi_{\bm{k}}\equiv\Pi^{+}_{\bm{k}}+\Pi^{-}_{\bm{k}}=\sigma(\epsilon_{p_{x},p_{z}}^{+}(p_{x},p_{z})+\epsilon_{p_{x},p_{z}}^{-}(p_{x},p_{z}))k_{f}/p$, from Eq. (9). This energy flux is transmitted to Sector A at $p^{*}$, where it provides a boundary condition for Eqs. (12). Indeed, the energy flux at $p^{*}$, $\Pi^{*}=\sigma(\epsilon_{p_{x},p_{z}}^{+}(p_{x},p_{z})+\epsilon_{p_{x},p_{z}}^{-}(p_{x},p_{z}))k_{f}/p^{*}$, determines the values of correlators $\langle vv\rangle_{\bm{p}\bm{k}}(p^{*})$, $\langle\eta v\rangle_{\bm{p}\bm{k}}(p^{*})$ and $\langle\eta\eta\rangle_{\bm{p}\bm{k}}(p^{*})$. Assuming zero helicity injection $\chi_{\bm{p}}^{s}=\chi_{\bm{p}}^{-s}$, system (71) is then solved by:
$\langle vv\rangle_{\bm{p}\bm{k}}$ $=\frac{\Pi^{*}}{\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}p^{*2}}{p^{4}}$ (13)
$\langle\eta v\rangle_{\bm{p}\bm{k}}$ $=i\frac{p_{z}}{p_{x}}\frac{1}{p^{2}}\int_{p_{y}^{*}}^{p_{y}}q^{2}\langle vv\rangle_{\bm{q}\bm{k}}~dq_{y}$ (14)
@@
$\langle\eta\eta\rangle_{\bm{p}\bm{k}}$ $=\frac{p_{z}}{p_{x}}\int_{p_{y}^{*}}^{p_{y}}2\Im\langle\eta v\rangle_{\bm{q}\bm{k}}dq_{y}+\frac{p_{0}^{2}\Pi^{*}}{\sqrt{2}|{U^{\prime}}p_{x}|}$ (16)
@@
with $p_{0}^{2}=p_{x}^{2}+p_{z}^{2}$ and $p_{y}^{*}=(p^{*2}-p_{0}^{2})^{\frac{1}{2}}$. From Eq. (17), we obtain a finite correlator $\langle\eta\eta\rangle_{\bm{p}\bm{k}}$ at large $p$, responsible for the creation of small-scale cross-flow enstrophy. The small-scale energy flux is then obtained from
$\Pi_{\bm{p}}=-\frac{p_{x}{U^{\prime}}}{\sqrt{2}(p_{x}^{2}+p_{z}^{2})}\Big(p^{2}\langle vv\rangle_{\bm{p}\bm{k}}+\langle\eta\eta\rangle_{\bm{p}\bm{k}}\Big)$
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}=\sqrt{2}{U^{\prime}}\left(-\frac{p_{x}p_{y}}{p_{x}^{2}+p_{z}^{2}}\langle vv\rangle_{\bm{p}\bm{k}}+\frac{p_{z}}{p_{x}^{2}+p_{z}^{2}}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}\right)$
@@
where $\epsilon_{p_{x},p_{z}}=2\epsilon^{s}_{p_{x},p_{z}}$ denotes the total energy injected in each line $(p_{x},p_{z})$. Figure 1(e) shows this energy transfer, obtained for modes with $p^{*}>k_{f}$. A similar approach can be used for slower modes $p^{*}<k_{f}$, see SM §E.2, and whose total contribution is found to be negligible. For $p>k_{f}$, the energy transfer $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ is positive in Sector H and negative in Sector A, where waves extract energy from the 2D flow.
@@
$\frac{1}{4Ro_{\epsilon}^{2}}\left(\frac{{U^{\prime}}}{\Omega}\right)^{2}=\frac{T_{\rm 3D-2D}}{\epsilon}\Big[\frac{{U^{\prime}}}{\Omega}\Big]=\frac{\mathcal{V}}{\sqrt{2}}\int_{\begin{subarray}{c}s=\pm 1\end{subarray}}d{\bm{p}}{U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}$ (20)
@@
$(ii)$ When $\tilde{s}=s$, the resonance condition (3) can be smoothened to better capture interactions when $\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{s}\sim{U^{\prime}}$, whose contribution is never fully zero nor one. This can be achieved by multiplying the Reynolds stress in (19) by a smooth Lorentzian filter [lvov1997statistical, gome2025waves],
$F_{k_{f}}=\left(1+\left(\frac{\Omega}{{U^{\prime}}}Q\frac{4\pi p_{z}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}{k_{f}^{3}}\right)^{2}\right)^{-1/2}$ (21)
In a similar vein, one can refine the resonant condition (3) when $\tilde{s}=-s$ by introducing a parameter $\tilde{\omega}$ in the resonant condition $\omega_{\bm{p}}^{+}+\omega_{\bm{q}}^{-}<\tilde{\omega}{U^{\prime}}$. This is equivalent to changing $\frac{{U^{\prime}}}{\Omega}\to\tilde{\omega}\frac{{U^{\prime}}}{\Omega}$ in the theory. Both refinements of (3) can be required to correct our approximate treatment of near-resonances (see [gome2025waves] and SM §B).
In Fig. 3, solid lines show a numerical solution of the QL system. To produce it, we first integrate numerically the Reynolds stress (given by Eq. (19) for large $p^{*}>k_{f}$, but written more generally in SM Eq. (113)) over a discrete set of wavenumbers $p_{z}=2\pi m_{z}/L_{z}$ and $p_{x}=2\pi m_{x}/L_{x}$, ($m_{z},m_{x}\in\mathbb{Z}^{*}$). We then determine the condensate amplitude ${U^{\prime}}/\Omega$ and the resulting transfer $T_{\rm 3D-2D}$ by solving Eq. 20 in variable ${U^{\prime}}/\Omega$ via a root-finding procedure. In this integration, both $\tilde{\omega}$ and $Q$ are tuning parameters. In Fig. 3, we use $\tilde{\omega}=1$ and $Q=8$. Note that a better agreement with the DNS can be achieved when taking $\tilde{\omega}=1.2$ (See SM §F.2).
Using the smooth filter (21) results in a smooth transition between the homochiral and the heterochiral-interaction regimes around $Ro_{\epsilon}\simeq 0.1$. Note, however, that this smooth filter is not relevant in the $Ro\to 0$ limit [gome2025waves]. Finally, while capturing the disappearance of condensates as $Ro\to 0.5$, the introduction of cutoff $k_{\Omega}$ causes overshoots before the decay in condensate amplitude. We discuss this effect and its impact on the $Re$-dependence of the QL solution in SM §F.2.
@@ Appendix A Numerical setup
$\partial_{t}\bm{u}+\bm{u}\cdot\nabla\bm{u}$ $=-2\Omega\mathbf{e}_{z}\times\bm{u}-\nabla p+\nu\nabla^{2}\bm{u}+\nu_{h}\nabla^{2h}\bm{u}+\bm{f},$ (22)
in a periodic domain $(L_{x},L_{y},L_{z})$ of renormalized volume $\mathcal{V}=(L_{x}L_{y}L_{z})/(2\pi)^{3}$. Forcing $\bm{f}$ is such that $\langle\bm{f}_{\bm{p}}(t)\bm{f}_{\bm{q}}(t^{\prime})\rangle=2\epsilon\chi_{\bm{p}}\delta_{\bm{p}+\bm{q}}\delta(t-t^{\prime})$, with correlation $\chi_{\bm{p}}=\delta_{p-k_{f}\pm 1}/(4\pi k_{f}^{2}dk\mathcal{V})$ and $\epsilon$ the energy injection rate. Nondimensional Reynolds and Rossby numbers at the forcing scale $l_{f}\equiv 2\pi/k_{f}$ are defined as $Re=\epsilon^{1/3}k_{f}^{-4/3}/\nu$ and $Ro=\epsilon^{1/3}k_{f}^{2/3}/(2\Omega)$. We simulate Eq. (22) with the pseudospectral code GHOST [mininni2011hybrid], with resolution $(N_{x},N_{y},N_{z})=128(L_{x},L_{y},L_{z})/\pi$. The inertial range of the forward energy cascade is controlled by 8-order hyperviscosity $\nu_{8}=6\times 10^{-29}$.
We decompose the 3D field into helical modes, $\bm{u}(\bm{x},t)=\sum_{\bm{p},s=\pm 1}\bm{h}_{\bm{p}}^{s}~a_{\bm{p}}^{s}(t)~e^{i\omega_{\bm{p}}^{s}t+i\bm{p}\cdot\bm{x}}$, with the helical basis
$\bm{h}_{\bm{p}}^{s}=\frac{\bm{p}\times[\bm{p}\times\mathbf{e}_{z}]-isp\bm{p}\times\mathbf{e}_{z}}{\sqrt{2}pp_{\perp}}=\frac{1}{\sqrt{2}pp_{\perp}}\begin{bmatrix}p_{x}p_{z}-ispp_{y}\\ p_{z}p_{y}+ispp_{x}\\ -p_{x}^{2}-p_{y}^{2}\end{bmatrix}$, (23)
with $\bm{p}=(p_{x},p_{y},p_{z})$, $p=|\bm{p}|$ and $p_{\perp}^{2}=p_{x}^{2}+p_{y}^{2}$ [lesieur1987turbulence]. This basis satisfies $i\bm{p}\times\bm{h}_{\bm{p}}^{s}=s\bm{h}_{\bm{p}}^{s}$ hence diagonalizes the curl operator. From such decomposition on the instantaneous field $\bm{u}$, we then measure the 3D-2D energy transfer
$\Pi^{s_{1},s_{2}}=-\sum_{|\bm{k}|<5}\sum_{\bm{k}\in 2D,\bm{p},\bm{q}\in 3D}\langle\bm{u}_{\bm{k}}\cdot(\bm{u}_{\bm{p}}^{s_{1}}\times{\bm{\omega}_{\bm{q}}^{{s_{2}}})}\rangle\delta_{\bm{k}\bm{p}\bm{q}},$ (24)
Figure 1(d) in the main text shows the energy transfer due to homochiral waves, $\Pi^{++}+\Pi^{--}$, and due to heterochiral waves, $\Pi^{+-}+\Pi^{-+}$, for various values of $Ro$ and $Re$. Note that we only decompose the energy flux with regard to the chiralities $s_{1},s_{2}$ of the 3D modes, and do not further decompose contributions from the two chiralities of the 2D modes, as done in [buzzicotti2018energy, biferale2012inverse]. Throughout this paper, we always refer to $\bm{k}$ as being a 2D mode, and $\bm{p}$ and $\bm{q}$ as 3D waves.
@@ Appendix B Mean-wave kinetic theory
We consider a stationary mean flow $\bm{U}=\langle\bm{u}\rangle=U(y)\mathbf{e}_{x}$ consisting of two $x$-invariant jets, corresponding to the condensate in the rectangular domain $L_{y}=2L_{x}$. The average $\langle\cdot\rangle$ denotes both ensemble and vertical average. In the following, we denote by
@@
Interactions with the condensate dominate over eddy-eddy or wave-wave interactions: $1/{U^{\prime}}\ll\tau_{nl}$, hence $\bm{u}\cdot\nabla\bm{u}\ll\bm{u}\cdot\nabla\bm{U}$. Under this quasi-linear (QL) approximation [marston2023recent], the momentum equations for the mean flow and fluctuations $\bm{u}^{\prime}$ are written as
@@
$\partial_{t}\bm{u}^{\prime}+U(y)\partial_{x}\bm{u}^{\prime}$ $=-\partial_{y}Uv\mathbf{e}_{x}-2\Omega\mathbf{e}_{z}\times\bm{u^{\prime}}-\nabla p^{\prime}+\bm{f},$ (27)
where $u$ and $v$ denote the $x$ and $y$ components of the fluctuating velocity $\bm{u}^{\prime}$.
We decompose the fluctuation field as $\bm{u}^{\prime}(\bm{x},t)=\sum_{\bm{p},s=\pm 1}\bm{h}_{\bm{p}}^{s}~a_{\bm{p}}^{s}(t)~e^{i\omega_{\bm{p}}^{s}t+i\bm{p}\cdot\bm{x}}$, where $a_{\bm{p}}^{s}(t)$ is the time-varying amplitude of a helical mode $(\bm{p},s)$ and $\omega_{\bm{p}}^{s}$ is the frequency of the corresponding inertial wave. We write the energy balance for the waves $e_{\bm{p}}^{s}=\frac{1}{2}\langle a_{\bm{p}}^{s}a_{\bm{p}}^{s*}\rangle$, as an energy increment at current time $t$, from an arbitrary $t=0$,
$\frac{e^{s}_{\bm{p}}(t)-e^{s}_{\bm{p}}(0)}{t}$ $=\frac{1}{2t}\sum_{\bm{q},\bm{k},\tilde{s}}\Big[V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}}U_{\bm{k}}^{*}~\langle a^{s*}_{\bm{p}}a_{\bm{q}}^{\tilde{s}*}\rangle(t)e^{-i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}\Big]\delta_{\bm{k}\bm{p}\bm{q}}+c.c+\epsilon\chi^{s}_{\bm{p}}$ (28)
@@
$\omega_{\bm{p}\bm{q}}^{s\tilde{s}}\equiv\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{-s}=s2\Omega p_{z}\left(\frac{1}{p}+\frac{1}{q}\right)$ (29)
@@
$\frac{e^{s}_{\bm{p}}(t)-e^{s}_{\bm{p}}(0)}{t}$ $=\frac{1}{2}\sum_{\bm{q},\bm{k},\tilde{s}}\Big[V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}}U_{\bm{k}}^{*}~\langle a^{s*}_{\bm{p}}a_{\bm{q}}^{\tilde{s}*}\rangle\frac{1}{t}\int_{0}^{t}e^{-i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t^{\prime}}dt^{\prime}\Big]\delta_{\bm{k}\bm{p}\bm{q}}+c.c+\epsilon\chi^{s}_{\bm{p}}$
$\approx\frac{1}{2}\sum_{\bm{q},\bm{k},\tilde{s}}\frac{1-e^{-i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}}{i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}\Big[V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}}U_{\bm{k}}^{*}~\langle a^{s*}_{\bm{p}}a_{\bm{q}}^{\tilde{s}*}\rangle\Big]\delta_{\bm{k}\bm{p}\bm{q}}+c.c+\epsilon\chi^{s}_{\bm{p}}$
$=\frac{1}{2}\sum_{\bm{q},\bm{k},\tilde{s}}\Delta^{s\tilde{s}}(t)\Big[V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}}U_{\bm{k}}^{*}~\langle a^{s*}_{\bm{p}}a_{\bm{q}}^{\tilde{s}*}\rangle\Big]\delta_{\bm{k}\bm{p}\bm{q}}+c.c+\epsilon\chi^{s}_{\bm{p}}+o\left(\frac{{U^{\prime}}}{\Omega}\right)$ (30)

### v2

@@ Acknowledgements.
For modes with large-enough $p_{z}$, such that $p^{*}=4|p_{z}|\Omega/{U^{\prime}}>k_{f}$, and for ${U^{\prime}}/\Omega\lesssim 1$, forcing occurs in Sector H and waves there generate a decaying forward energy flux, $\Pi_{\boldsymbol{k}}\equiv\Pi^{+}_{\boldsymbol{k}}+\Pi^{-}_{\boldsymbol{k}}=\sigma(\epsilon_{p_{x},p_{z}}^{+}(p_{x},p_{z})+\epsilon_{p_{x},p_{z}}^{-}(p_{x},p_{z}))k_{f}/p$, from Eq. (9). This energy flux is transmitted to Sector A at $p^{*}$, where it provides a boundary condition for Eqs. (12). Indeed, the energy flux at $p^{*}$, $\Pi^{*}=\sigma(\epsilon_{p_{x},p_{z}}^{+}(p_{x},p_{z})+\epsilon_{p_{x},p_{z}}^{-}(p_{x},p_{z}))k_{f}/p^{*}$, determines the values of correlators $\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}(p^{*})$, $\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}(p^{*})$ and $\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}(p^{*})$. Assuming zero helicity injection $\chi_{\boldsymbol{p}}^{s}=\chi_{\boldsymbol{p}}^{-s}$, system (71) is then solved by:
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\frac{\Pi^{*}}{\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}p^{*2}}{p^{4}}$ (13)
$\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=i\frac{p_{z}}{p_{x}}\frac{1}{p^{2}}\int_{p_{y}^{*}}^{p_{y}}q^{2}\langle vv\rangle_{\boldsymbol{q}\boldsymbol{k}}~dq_{y}$ (14)
@@
$\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\frac{p_{z}}{p_{x}}\int_{p_{y}^{*}}^{p_{y}}2\Im\langle\eta v\rangle_{\boldsymbol{q}\boldsymbol{k}}dq_{y}+\frac{p_{0}^{2}\Pi^{*}}{\sqrt{2}|{U^{\prime}}p_{x}|}$ (16)
@@
with $p_{0}^{2}=p_{x}^{2}+p_{z}^{2}$ and $p_{y}^{*}=(p^{*2}-p_{0}^{2})^{\frac{1}{2}}$. From Eq. (17), we obtain a finite correlator $\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}$ at large $p$, responsible for the creation of small-scale cross-flow enstrophy. The small-scale energy flux is then obtained from
$\Pi_{\boldsymbol{p}}=-\frac{p_{x}{U^{\prime}}}{\sqrt{2}(p_{x}^{2}+p_{z}^{2})}\Big(p^{2}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}\Big)$
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}=\sqrt{2}{U^{\prime}}\left(-\frac{p_{x}p_{y}}{p_{x}^{2}+p_{z}^{2}}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\frac{p_{z}}{p_{x}^{2}+p_{z}^{2}}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\right)$
@@
where $\epsilon_{p_{x},p_{z}}=2\epsilon^{s}_{p_{x},p_{z}}$ denotes the total energy injected in each line $(p_{x},p_{z})$. Figure 1(e) shows this energy transfer, obtained for modes with $p^{*}>k_{f}$. A similar approach can be used for slower modes $p^{*}<k_{f}$, see SM §E.2, and whose total contribution is found to be negligible. For $p>k_{f}$, the energy transfer $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ is positive in Sector H and negative in Sector A, where waves extract energy from the 2D flow.
@@
$\frac{1}{4Ro_{\epsilon}^{2}}\left(\frac{{U^{\prime}}}{\Omega}\right)^{2}=\frac{T_{\rm 3D-2D}}{\epsilon}\Big[\frac{{U^{\prime}}}{\Omega}\Big]=\frac{\mathcal{V}}{\sqrt{2}}\int_{\begin{subarray}{c}s=\pm 1\end{subarray}}d{\boldsymbol{p}}{U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ (20)
@@
$(ii)$ When $\tilde{s}=s$, the resonance condition (3) can be smoothened to better capture interactions when $\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{s}\sim{U^{\prime}}$, whose contribution is never fully zero nor one. This can be achieved by multiplying the Reynolds stress in (19) by a smooth Lorentzian filter [31, 26],
$F_{k_{f}}=\left(1+\left(\frac{\Omega}{{U^{\prime}}}\frac{4\pi p_{z}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}{L_{y}k_{f}^{3}}\right)^{2}\right)^{-1}$ (21)
In a similar vein, one can refine the resonant condition (3) when $\tilde{s}=-s$ by introducing a parameter $\tilde{\omega}$ in the resonant condition $\omega_{\boldsymbol{p}}^{+}+\omega_{\boldsymbol{q}}^{-}<\tilde{\omega}{U^{\prime}}$. This is equivalent to changing $\frac{{U^{\prime}}}{\Omega}\to\tilde{\omega}\frac{{U^{\prime}}}{\Omega}$ in the theory. Both refinements of (3) can be required to correct our approximate treatment of near-resonances (see [26] and SM §B).
In Fig. 3, solid lines show a numerical solution of the QL system. To produce it, we first integrate numerically the Reynolds stress (given by Eq. (19) for large $p^{*}>k_{f}$, but written more generally in SM Eq. (113)) over a discrete set of wavenumbers $p_{z}=2\pi m_{z}/L_{z}$ and $p_{x}=2\pi m_{x}/L_{x}$, ($m_{z},m_{x}\in\mathbb{Z}^{*}$). We then determine the condensate amplitude ${U^{\prime}}/\Omega$ and the resulting transfer $T_{\rm 3D-2D}$ by solving Eq. 20 in variable ${U^{\prime}}/\Omega$ via a root-finding procedure. In this integration, $\tilde{\omega}$ is a tuning parameter. In Fig. 3, we use $\tilde{\omega}=1$. Note that a better agreement with the DNS can be achieved when taking $\tilde{\omega}=1.2$ (See SM §F.2).
Using the smooth filter (21) results in a smooth transition between the homochiral and the heterochiral-interaction regimes around $Ro_{\epsilon}\simeq 0.1$. Note, however, that this smooth filter is not relevant in the $Ro\to 0$ limit [26]. Finally, while capturing the disappearance of condensates as $Ro\to 0.5$, the introduction of cutoff $k_{\Omega}$ causes overshoots before the decay in condensate amplitude. We discuss this effect and its impact on the $Re$-dependence of the QL solution in SM §F.2.
@@ Appendix A Numerical setup
$\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}$ $=-2\Omega\mathbf{e}_{z}\times\boldsymbol{u}-\nabla p+\nu\nabla^{2}\boldsymbol{u}+\nu_{h}\nabla^{2h}\boldsymbol{u}+\boldsymbol{f},$ (22)
in a periodic domain $(L_{x},L_{y},L_{z})$ of renormalized volume $\mathcal{V}=(L_{x}L_{y}L_{z})/(2\pi)^{3}$. Forcing $\boldsymbol{f}$ is such that $\langle\boldsymbol{f}_{\boldsymbol{p}}(t)\boldsymbol{f}_{\boldsymbol{q}}(t^{\prime})\rangle=2\epsilon\chi_{\boldsymbol{p}}\delta_{\boldsymbol{p}+\boldsymbol{q}}\delta(t-t^{\prime})$, with correlation $\chi_{\boldsymbol{p}}=\delta_{p-k_{f}\pm 1}/(4\pi k_{f}^{2}dk\mathcal{V})$ and $\epsilon$ the energy injection rate. Nondimensional Reynolds and Rossby numbers at the forcing scale $l_{f}\equiv 2\pi/k_{f}$ are defined as $Re=\epsilon^{1/3}k_{f}^{-4/3}/\nu$ and $Ro=\epsilon^{1/3}k_{f}^{2/3}/(2\Omega)$. We simulate Eq. (22) with the pseudospectral code GHOST [38], with resolution $(N_{x},N_{y},N_{z})=128(L_{x},L_{y},L_{z})/\pi$. The inertial range of the forward energy cascade is controlled by 8-order hyperviscosity $\nu_{8}=6\times 10^{-29}$.
We decompose the 3D field into helical modes, $\boldsymbol{u}(\boldsymbol{x},t)=\sum_{\boldsymbol{p},s=\pm 1}\boldsymbol{h}_{\boldsymbol{p}}^{s}~a_{\boldsymbol{p}}^{s}(t)~e^{i\omega_{\boldsymbol{p}}^{s}t+i\boldsymbol{p}\cdot\bm{x}}$, with the helical basis
$\boldsymbol{h}_{\boldsymbol{p}}^{s}=\frac{\boldsymbol{p}\times[\boldsymbol{p}\times\mathbf{e}_{z}]-isp\boldsymbol{p}\times\mathbf{e}_{z}}{\sqrt{2}pp_{\perp}}=\frac{1}{\sqrt{2}pp_{\perp}}\begin{bmatrix}p_{x}p_{z}-ispp_{y}\\ p_{z}p_{y}+ispp_{x}\\ -p_{x}^{2}-p_{y}^{2}\end{bmatrix}$, (23)
with $\boldsymbol{p}=(p_{x},p_{y},p_{z})$, $p=|\boldsymbol{p}|$ and $p_{\perp}^{2}=p_{x}^{2}+p_{y}^{2}$ [34]. This basis satisfies $i\boldsymbol{p}\times\boldsymbol{h}_{\boldsymbol{p}}^{s}=s\boldsymbol{h}_{\boldsymbol{p}}^{s}$ hence diagonalizes the curl operator. From such decomposition on the instantaneous field $\boldsymbol{u}$, we then measure the 3D-2D energy transfer
$\Pi^{s_{1},s_{2}}=-\sum_{|\boldsymbol{k}|<5}\sum_{\boldsymbol{k}\in 2D,\boldsymbol{p},\boldsymbol{q}\in 3D}\langle\boldsymbol{u}_{\boldsymbol{k}}\cdot(\boldsymbol{u}_{\boldsymbol{p}}^{s_{1}}\times{\boldsymbol{\omega}_{\boldsymbol{q}}^{{s_{2}}})}\rangle\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}},$ (24)
Figure 1(d) in the main text shows the energy transfer due to homochiral waves, $\Pi^{++}+\Pi^{--}$, and due to heterochiral waves, $\Pi^{+-}+\Pi^{-+}$, for various values of $Ro$ and $Re$. Note that we only decompose the energy flux with regard to the chiralities $s_{1},s_{2}$ of the 3D modes, and do not further decompose contributions from the two chiralities of the 2D modes, as done in [10, 7]. Throughout this paper, we always refer to $\boldsymbol{k}$ as being a 2D mode, and $\boldsymbol{p}$ and $\boldsymbol{q}$ as 3D waves.
@@ Appendix B Mean-wave kinetic theory
We consider a stationary mean flow $\boldsymbol{U}=\langle\boldsymbol{u}\rangle=U(y)\mathbf{e}_{x}$ consisting of two $x$-invariant jets, corresponding to the condensate in the rectangular domain $L_{y}=2L_{x}$. The average $\langle\cdot\rangle$ denotes both ensemble and vertical average. In the following, we denote by
@@
Interactions with the condensate dominate over eddy-eddy or wave-wave interactions: $1/{U^{\prime}}\ll\tau_{nl}$, hence $\boldsymbol{u}\cdot\nabla\boldsymbol{u}\ll\boldsymbol{u}\cdot\nabla\boldsymbol{U}$. Under this quasi-linear (QL) approximation [37], the momentum equations for the mean flow and fluctuations $\boldsymbol{u}^{\prime}$ are written as
@@
$\partial_{t}\boldsymbol{u}^{\prime}+U(y)\partial_{x}\boldsymbol{u}^{\prime}$ $=-\partial_{y}Uv\mathbf{e}_{x}-2\Omega\mathbf{e}_{z}\times\boldsymbol{u^{\prime}}-\nabla p^{\prime}+\boldsymbol{f},$ (27)
where $u$ and $v$ denote the $x$ and $y$ components of the fluctuating velocity $\boldsymbol{u}^{\prime}$.
We decompose the fluctuation field as $\boldsymbol{u}^{\prime}(\boldsymbol{x},t)=\sum_{\boldsymbol{p},s=\pm 1}\boldsymbol{h}_{\boldsymbol{p}}^{s}~a_{\boldsymbol{p}}^{s}(t)~e^{i\omega_{\boldsymbol{p}}^{s}t+i\boldsymbol{p}\cdot\bm{x}}$, where $a_{\boldsymbol{p}}^{s}(t)$ is the time-varying amplitude of a helical mode $(\boldsymbol{p},s)$ and $\omega_{\boldsymbol{p}}^{s}$ is the frequency of the corresponding inertial wave. We write the energy balance for the waves $e_{\boldsymbol{p}}^{s}=\frac{1}{2}\langle a_{\boldsymbol{p}}^{s}a_{\boldsymbol{p}}^{s*}\rangle$, as an energy increment at current time $t$, from an arbitrary $t=0$,
$\frac{e^{s}_{\boldsymbol{p}}(t)-e^{s}_{\boldsymbol{p}}(0)}{t}$ $=\frac{1}{2t}\sum_{\boldsymbol{q},\boldsymbol{k},\tilde{s}}\Big[V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}}U_{\boldsymbol{k}}^{*}~\langle a^{s*}_{\boldsymbol{p}}a_{\boldsymbol{q}}^{\tilde{s}*}\rangle(t)e^{-i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}\Big]\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}+c.c+\epsilon\chi^{s}_{\boldsymbol{p}}$ (28)
@@
$\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}\equiv\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{-s}=s2\Omega p_{z}\left(\frac{1}{p}+\frac{1}{q}\right)$ (29)
@@
$\frac{e^{s}_{\boldsymbol{p}}(t)-e^{s}_{\boldsymbol{p}}(0)}{t}$ $=\frac{1}{2}\sum_{\boldsymbol{q},\boldsymbol{k},\tilde{s}}\Big[V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}}U_{\boldsymbol{k}}^{*}~\langle a^{s*}_{\boldsymbol{p}}a_{\boldsymbol{q}}^{\tilde{s}*}\rangle\frac{1}{t}\int_{0}^{t}e^{-i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t^{\prime}}dt^{\prime}\Big]\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}+c.c+\epsilon\chi^{s}_{\boldsymbol{p}}$
$\approx\frac{1}{2}\sum_{\boldsymbol{q},\boldsymbol{k},\tilde{s}}\frac{1-e^{-i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}}{i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}\Big[V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}}U_{\boldsymbol{k}}^{*}~\langle a^{s*}_{\boldsymbol{p}}a_{\boldsymbol{q}}^{\tilde{s}*}\rangle\Big]\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}+c.c+\epsilon\chi^{s}_{\boldsymbol{p}}$
$=\frac{1}{2}\sum_{\boldsymbol{q},\boldsymbol{k},\tilde{s}}\Delta^{s\tilde{s}}(t)\Big[V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}}U_{\boldsymbol{k}}^{*}~\langle a^{s*}_{\boldsymbol{p}}a_{\boldsymbol{q}}^{\tilde{s}*}\rangle\Big]\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}+c.c+\epsilon\chi^{s}_{\boldsymbol{p}}+o\left(\frac{{U^{\prime}}}{\Omega}\right)$ (30)